An engineer wants to make a 40% cement mixture. How many pounds of a 60% cement mixture must she mix with 12 pounds of a 30% mixture to create a 40% cement mixture?

Answer by stanbon(75887)

An engineer wants to make a 40% cement mixture. How many pounds of a 60% cement mixture must she mix with 12 pounds of a 30% mixture to create a 40% cement mixture?
------
Equation:
cement + cement = cement
0.60x + 0.30*12 = 0.40(x+12)
---
Multiply thru by 100 to get:
60x + 30*12 = 40x + 40*12
20x = 10*12
x = 6 lbs. (amt. of 60% mixture needed)
================
Cheers,
Stan H.
==================